Find the derivative, r(t), of the vector function. r(t) = eti- j+ln(1 + 7t)k r(t) =Find parametric equations for the tangent line to the curve with the given parametric equations at the specified point. T = 5
Answer #1

soli 8(t)= e inj+ ln (147t) i 7 8 7 to 8(t) = 8 t ei-o + IC 7t+1 8 7k 6/8 (t) = 8 ttet i + Oj+ 7t+1 Solin 16=56t; g = t et, z(5,0,0) = (0,, Y, Z,) 25,1,17 = (a,b,c) z) Lect) = 5 +st a) y (t)= ott 27 Elt) Ott :sse (t), y(t), z1t) +(5+5+, t, t)